### v1

@@ 3.1 Parallel Convolutions (PC)
This approach contrasts with the inception block [30], which typically includes multiple parallel paths starting from the same input, each with different combinations of convolutions and sometimes pooling operations, without a shared embedded convolution. The inception block aims to capture multi-scale information by applying various-sized convolutions (e.g., , , ) in parallel and then merging their outputs. However, each path in an inception block operates independently from the start, without a shared feature extraction stage. This increases the number of parameters and may lead to redundancy, as each path may learn similar features.
@@ 3.2 SE Attention Block
Based on rigorous evaluation of attention mechanisms in Section 5, we incorporate an SE block [22] into our final model. This block is recognized for its efficiency and lightweight nature. The SE block is composed of two distinct operations: Squeeze and Excitation. In the squeeze phase, input images of size are transformed to a format through a Global Average Pooling (GAP) layer, which compresses the spatial resolutions, retaining only channel-centric information for the subsequent excitation operation. The excitation phase employs a series of layers, beginning with a fully connected layer complemented by a reduction factor . This is then subjected to ReLU activation, succeeded by another fully connected layer, culminating in a sigmoid activation to produce the final output of the Excitation operation. A scaling transformation is executed to assimilate the channel-specific data, yielding an output enriched with channel-level information.
@@ 4.1 Data and Pre-processing
The proposed model is trained and validated using the Brain Tumor Segmentation (BraTS) benchmark datasets: BraTS 2020 [1, 31, 32] and BraTS 2021 [1]. BraTS 2021, is a superset of BraTS 2020, encompasses patients, which include high-grade gliomas (HGG) and low-grade gliomas (LGG). The BraTS 2020 dataset contains patients, of which have been diagnosed with LGG, with the remainder having HGG.
The BraTS 2020 and BraTS 2021 datasets are used for efficiency and sustainability, with initial testing on the smaller dataset allowing for algorithm refinement and model selection before deployment on the larger dataset. It conserves computational resources and time, and aligns with iterative development best practices, where initial testing on a subset of data can provide quick feedback, critical for early-stage model tuning and optimization [33].
Each dataset consists of 3D scans with individual “slices” or images, each having pixels. These scans capture four MRI modalities–T2, T1, T1ce, and FLAIR–crucial for brain tumor segmentation, offering distinct insights. T1-weighted images illustrate anatomical structures, distinguishing between gray and white matter. T2-weighted images aid in visualizing edema, by emphasizing water content. T1ce images, with contrast enhancement, focus on blood vessel imaging, a key component in identifying active tumor growth. Conversely, FLAIR images suppress cerebrospinal fluid signals, illuminating anomalies in intensity and subtle lesions indicative of tumor expansion. Ground truth segmentation masks are meticulously annotated by one to four expert neuroradiologists per case. The scan has been segmented into four primary classes: background (BG, Label 0), necrotic and non-enhancing tumor (NCR/NET, Label 1), edema (ED, Label 2), and enhancing tumor (ET, Label 4) as exemplified in Figure 3. Following common practices in the literature, we classify these into three main tumor regions for segmentation: whole tumor (WT) encompassing NCR/NET, ED, and ET (Labels 1, 2, 4), tumor core (TC) consisting of NCR/NET and ET (Labels 1, 4), and the sole ET (Label 4).
@@ 4.2 Implementation Details
We implement our network via Keras in Tensorflow 2.15. Computations are executed on a single NVIDIA V100 16GB GPU. However, it’s important to note that our architecture is versatile and can be run on other graphics cards, including consumer-grade GPUs. For training, we employ the ADAM optimizer [35], setting the learning rate to . Training proceeds with a batch size of , a choice primarily dictated by GPU memory constraints. We use a constant number of epochs. To mitigate the risk of overfitting and enhance the model’s generalization capabilities, L2 regularization is applied to the convolutional kernel parameters with a factor of .
During preliminary experimentation, we evaluated the ReduceLROnPlateau callback for dynamic learning rate adjustments. However, observations indicated a predisposition towards overfitting when employed. As such, it was excluded from final training (see Section 5.1.1).
The specific hyper-parameter settings we adopted during model training are detailed in Table 1. Detailed results with a discussion are in Section 5. The source code is available at [36] with final models and analysis results at [37].
@@ 4.3 Loss Function
where , is the voxel count for class , and is the total count of voxels across all classes. This formula ensures that classes with fewer voxels receive higher weights to balance the loss during the training process. Here we use the WT, TC and ET regions instead of the individual output channels above to compute the weights and we get , , and .
Overall this givens our Weighted Dice score Loss (WDL),
@@
It is also important to note that while the network output channels include the background class (BG), it is not included in the loss and weights calculation. We found its presence improves the segmentation performance of the model. We also explored the use of different output channels, weights, manual adjustment of the weights. However, it did not yield satisfactory results.
@@ 4.4 Evaluation Metrics
We measure the effectiveness of the proposed model using the Dice similarity coefficient (DSC), and the percentile Hausdorff distance (HD95). DSC and HD95 are widely accepted as the primary performance evaluation metrics in image segmentation tasks. The DSC quantifies the spatial overlap between the ground truth and the predicted segmentation region. HD95 calculates the percentile of the distances between the points in the ground truth and the predicted set. This is akin to the conventional symmetric Hausdorff distance but reduces the impact of outliers by focusing on the percentile. HD95 in particular indicates the accuracy of boundary prediction, and a lower score reveals the model’s precision in delineating the tumor margins.
@@
We employ different metrics for evaluating our model based on the data partitioning approach used. For model selection (see Section 5.1, and Section 5.2), we employ an / training-testing holdout split, and we report the mean and standard deviation of the per-sample (per-patient) metrics to gain insights into the model’s performance under controlled conditions. Once the optimal model (LATUP-Net) is determined, and for the comparison to other state-of-the-arts models (see Section 5.4), five-fold cross-validation is applied to consider any data dependencies and ensure a comprehensive evaluation of the model’s performance across varied data scenarios. During cross-validation, we calculate the mean and standard deviation of the mean DSC and HD95 across all folds to assess the model’s robustness and general performance.
@@ 5.1 Overall Performance Analysis
The models are initially selected based on a single / split using the same training hyperparameters (see Table 1) on the BraTS 2020 dataset. Table 2 shows the segmentation results of the test set from these training runs. The table also lists the parameter counts, allowing for a comparison of model complexity. The performance of these models is assessed by employing two key metrics: per-sample DSC and HD95.
rowsep=3pt Model DSC () HD95 (mm) Parameters WT TC ET WT TC ET U-Net M Inception M PC M PC + SE M PC + WDL M PC + SE + WDL M
In the inception model, we modify the baseline U-Net by adding the inception module to its first block. We observe that the number of parameters decreased by M. The model outperforms the U-Net with a DSC increment by , , and and a decrement in the HD95 by , , and for WT, TC, and ET respectively. However, despite the reduction in the parameter counts, using the inception block led to the model requiring a lot of memory, thereby increasing its memory consumption during training. This memory-intensive behavior of the model might attributed to the inception block design.
The PC model has M fewer parameters than U-Net and exhibits rapid convergence during the initial training epochs, leading to shorter training time and reduced need for computational resources. It also obtains a significant improvement in the segmentation performance compared to U-Net with a DSC increment by , , and and HD95 decrement by , , and for WT, TC and ET respectively.
When comparing our proposed parallel convolutions (PC) with the inception module, PC offers a strategic advantage. By replacing the conventional convolution with a convolution, PC achieves a more spatially compact and contextually rich representation, while effectively reducing memory consumption. In contrast, the inception module with M more parameters faced a surge in model memory usage during training. Furthermore, our design’s judicious positioning of pooling operations optimally condenses feature map dimensions, ensuring efficient memory usage without compromising capturing features. The PC model outperformed the inception model with , , and DSC increment and , , and HD95 decrement for WT, TC, and ET respectively which proves PC’s ability to segment difficult tumor regions such as TC.
Integrating the attention mechanism, specifically SE attention with the PC model, leads to an increment in the model parameters but slightly improves the performance on the test data. It enhances the model’s ability to segment small regions, as evidenced by and increments in the DSC for WT and ET, respectively. In contrast, the DSC performance of TC detection decreased slightly by . At the same time, the benefits of the parallel convolutions are maintained, despite a slight trade-off for certain tumor regions. Several lightweight attention modules are compared in Section 5.2.
@@
This improvement underscores the utility of the WDL in case of a segmentation region size imbalance. However, we notice that when adding the attention mechanism, the result of WT and TC decreased slightly in both DCS and HD95, which leads us to check whether attention is needed. This is further investigated in Section 5.3.
Figure 4 depicts qualitative comparisons between the various networks in segmenting the distinct tumor regions. The qualitative results demonstrate that our LATUP-Net which consists of PC and SE trained with WDL, outperformed all other models, consistent with the quantitative results in Table 2.
@@ 5.2 Comparison of Different Attention Mechanisms
In our approach so far, we have used the SE attention mechanism. In this section, we explore whether alternative attention mechanisms (see Section 2.3) may yield improved performance within the confines of our model. For a uniform evaluation, we incorporate all attention mechanisms after the parallel convolutions (PC). It is also important to note that the model trained using dice loss and the results presented here are derived from configurations that do not include the weighted Dice score loss (WDL). This choice was made to isolate the impact of the attention mechanisms themselves, allowing for a clear comparison of their effectiveness in the absence of the WDL’s influence.
@@
Table 3 summarizes our findings, detailing the performance metrics of residual SE, CBAM, ECA, and their combinations. While all mechanisms enhance the model’s accuracy, SE stands out as the most effective.
CBAM, despite its comprehensive design, might be limited by its complexity, potentially impacting efficiency. Similarly, while ECA aims for computational efficiency, it might fall short in capturing intricate inter-channel dynamics as effectively as SE. SE’s distinct capability in identifying enhanced tumors, along with its reduced parameters, underscores its utility. Our results affirm the significance of SE in tumor detection but also suggest that the performance of other attention mechanisms could be context-dependent, varying with the dataset or task.
Analyzing the standard deviation values, the variance in DSC scores among methods appears minimal. Given this minimal divergence, the definitive superiority of certain attention mechanisms can be debated. In such scenarios, streamlined models with fewer parameters, such as PC+SE with M parameters, offer competitive or superior performance.
The HD95 metric presents more noticeable differences across methodologies. This metric sheds light on spatial discrepancies in segmentation outcomes, possibly arising from the mechanisms’ ability to capture spatial nuances, outliers, irregular tumor shapes, or unique dataset attributes.
rowsep=3pt Method DSC () HD95 (mm) Parameters WT TC ET WT TC ET PC+SE M PC+CBAM M PC+CBAM+SE M PC+ (SE-3D) M PC+Residual SE M PC+ECA M

### v2

@@ 3.1 Parallel Convolutions (PC)
This approach contrasts with the inception block [38], which typically includes multiple parallel paths starting from the same input, each with different combinations of convolutions and sometimes pooling operations, without a shared embedded convolution. The inception block aims to capture multi-scale information by applying various-sized convolutions (e.g., , , ) in parallel and then merging their outputs. However, each path in an inception block operates independently of the others, without a shared feature extraction stage. This increases the number of parameters and may lead to redundancy, as each path may learn similar features.
@@ 3.2 SE Attention Block
The attention mechanisms explored in this study include Squeeze-and-Excitation (SE) [19], the Convolutional Block Attention Module (CBAM) [18], Efficient Channel Attention (ECA) [39], and Residual Squeeze-and-Excitation (RSE) [40]. We further introduce a modified variant of SE where the fully connected (dense) layers are replaced by a 3D convolutional layer. We also experiment with combined mechanisms such as fusing CBAM and SE. The motivation behind combining CBAM and SE is to leverage the strengths of both. CBAM’s ability to focus on pertinent spatial and channel features and SE’s capacity to recalibrate channel-wise features may enhance the model’s ability to capture complex interdependencies in the data. Another approach is designed to exploit the strengths of convolutions while adhering to the principles of the SE mechanism. Replacing the dense layer in SE with a 3D convolution layer (SE-3D) is aimed at maintaining the spatial information of the input tensor and capturing local spatial correlations, while simultaneously maintaining the ability to recalibrate channel-wise features.
Based on rigorous evaluation of attention mechanisms in Section 5, we incorporate an SE block [19] into our final model, which is illustrated in Figure 3. This block is recognized for its efficiency and lightweight nature. The SE block is composed of two distinct operations: Squeeze and Excitation. In the squeeze phase, input images of size are transformed to a format through a Global Average Pooling (GAP) layer, which compresses the spatial resolutions, retaining only channel-centric information for the subsequent excitation operation. The excitation phase employs a series of layers, beginning with a fully connected layer complemented by a reduction factor . This is then subjected to ReLU activation, succeeded by another fully connected layer, culminating in a sigmoid activation to produce the final output of the Excitation operation. A scaling transformation is executed to assimilate the channel-specific data, yielding an output enriched with channel-level information.
@@ 4.1 Data and Pre-processing
The proposed model is trained and validated using the Brain Tumor Segmentation (BraTS) benchmark datasets: BraTS 2020 [1, 3, 4] and BraTS 2021 [1]. BraTS 2021, a superset of BraTS 2020, encompasses patients, including high-grade gliomas (HGG) and low-grade gliomas (LGG). The BraTS 2020 dataset contains patients, of which have been diagnosed with LGG, with the remainder having HGG.
The BraTS 2020 and BraTS 2021 datasets are used for efficiency and sustainability, with initial testing on the smaller dataset allowing for algorithm refinement and model selection before deployment on the larger dataset. It conserves computational resources and time, and aligns with iterative development best practices, where initial testing on a subset of the data can provide quick feedback, critical for early-stage model tuning and optimization [41].
Each dataset consists of 3D scans with individual “slices” or images, each having pixels. These scans capture four MRI modalities—T2, T1, T1ce, and FLAIR—crucial for brain tumor segmentation, offering distinct insights. T1-weighted images illustrate anatomical structures, distinguishing between gray and white matter. T2-weighted images aid in visualizing edema, by emphasizing water content. T1ce images, with contrast enhancement, focus on blood vessel imaging, a key component in identifying active tumor growth. Conversely, FLAIR images suppress cerebrospinal fluid signals, illuminating anomalies in intensity and subtle lesions indicative of tumor expansion. Ground truth segmentation masks are meticulously annotated by one to four expert neuroradiologists per case. The scan has been segmented into four primary classes: background (BG, Label 0), necrotic and non-enhancing tumor (NCR/NET, Label 1), edema (ED, Label 2), and enhancing tumor (ET, Label 4) as exemplified in Figure 4. Following common practices in the literature, we classify these into three main tumor regions for segmentation: whole tumor (WT) encompassing NCR/NET, ED, and ET (Labels 1, 2, 4), tumor core (TC) consisting of NCR/NET and ET (Labels 1, 4), and the sole ET (Label 4).
@@ 4.2 Implementation Details
We implemented our network via Keras in Tensorflow 2.15. Computations are executed on a single NVIDIA GeForce RTX3060 12 GB GPU, which is considered a relatively low-end consumer card. For training, we employed the ADAM optimizer [43], setting the learning rate to . Training proceeds with a batch size of , a choice primarily dictated by GPU memory constraints. We used a constant number of epochs. To mitigate the risk of overfitting and enhance the model’s generalization capabilities, L2 regularization was applied to the convolutional kernel parameters with a factor of . This regularization factor was selected based on experimentation during the model selection process, as detailed in the supplementary material.
Leaky ReLU with a leak factor of was used as the activation function for the hidden layers. This value is widely used in deep learning and was chosen based on its effectiveness in similar tasks, such as image segmentation, where it helps regularize the model by allowing gradient flow for negative inputs, contributing to stable training. Studies, including nnU-Net [44] and the original leaky ReLU paper [45], have shown that works well in practice.
During preliminary experimentation, we evaluated the ReduceLROnPlateau callback for dynamic learning rate adjustments. However, observations indicated a predisposition towards overfitting when it was employed. As such, it was excluded from the final training (see Section 5.1.1).
The specific hyper-parameter settings we adopted during model training are detailed in Table 2. Detailed results with a discussion are in Section 5 with additional information about the parameter selection process available in the supplementary material. The source code is available at [35] with final models and analysis results at [46].
@@ 4.3 Loss Function
where , is the voxel count for class , and is the total count of voxels across all classes. This formula ensures that classes with fewer voxels receive higher weights to balance the loss during the training process. Here we use the WT, TC and ET regions instead of the individual output channels above to compute the weights, and we get , , and .
Overall this gives our Weighted Dice score Loss (WDL),
@@
It is also important to note that while the network output channels include the background class (BG), it is not included in the loss and weights calculation. We found its presence improves the segmentation performance of the model. We also explored the use of different output channels, weights, and manual adjustment of the weights. However, it did not yield satisfactory results.
@@ 4.4 Evaluation Metrics
We measure the effectiveness of the proposed model using the Dice similarity coefficient (DSC), and the percentile Hausdorff distance (HD95). DSC and HD95 are widely accepted as the primary performance evaluation metrics in image segmentation tasks. The DSC quantifies the spatial overlap between the ground truth and the predicted segmentation region. HD95 calculates the percentile of the distances between the points in the ground truth and the predicted set. This is akin to the conventional symmetric Hausdorff distance but reduces the impact of outliers by focusing on the percentile. HD95 in particular indicates the accuracy of boundary prediction, revealing the model’s precision in delineating the tumor margins.
@@
We employ different metrics for evaluating our model based on the data partitioning approach used. For model selection (see Section 5.1, and Section 5.2), we employ an / training-testing holdout split, and we report the mean and standard deviation of the per-sample (per-patient) metrics to gain insights into the model’s performance under controlled conditions. Once the optimal model (LATUP-Net) is determined, and for comparison to the state-of-the-arts models (see Section 5.5), five-fold cross-validation is applied to consider any data dependencies and ensure a comprehensive evaluation of the model’s performance across varied data scenarios. During cross-validation, we calculate the mean and standard deviation of the mean DSC and HD95 across all folds to assess the model’s robustness and general performance.
@@ 5.1 Overall Performance Analysis
The models are initially selected based on a single / split using the same training hyperparameters (see Table 2) on the BraTS 2020 dataset. Table 5.1 shows the segmentation results of the test set from these training runs. The performance is assessed by employing two key metrics: per-sample DSC and HD95.
rowsep=3pt
{tblr}
colspec=Q[2.5cm,valign=m]lllllllll,
cell1-21-8=gray!20,
cell11=r=2l,
cell12=c=3l,
cell15=c=3l,
cell18=r=2l
Model DSC () HD95 (mm)
WT TC ET WT TC ET
U-Net
Inception
PC
PC + SE
PC + WDL
PC + SE + WDL
In the Inception model, we replace the first U-Net block with an inception module. This modification improves the segmentation results compared to the U-Net, with DSC improvements of , , and for whole tumor (WT), tumor core (TC), and enhancing tumor (ET), while reducing the HD95 by , , and for WT, TC, and ET, respectively. However, despite these performance gains, the Inception model significantly increases memory usage during training, likely due to the module’s more complex structure, which combines multiple convolutional and pooling operations.
The PC model, using parallel convolutions, demonstrates superior efficiency compared to both U-Net and Inception. It achieves faster convergence during training and reduces the need for computational resources. Additionally, PC provides substantial improvements in segmentation performance compared to U-Net, with DSC gains of , , and and HD95 reductions of , , and for WT, TC, and ET, respectively.
When comparing our proposed parallel convolutions (PC) with the Inception module, PC offers a strategic advantage. By replacing the conventional convolution with a convolution, PC achieves a more spatially compact and contextually rich representation, while effectively reducing memory consumption. In contrast, the Inception module with its complex configuration faced a surge in model memory usage during training. Furthermore, our design’s judicious positioning of pooling operations optimally condenses feature map dimensions, ensuring efficient memory usage without compromising capturing features. The PC model outperforms the inception model with , , and DSC increment and , , and HD95 decrement for WT, TC, and ET, respectively, which proves PC’s ability to segment difficult tumor regions such as TC.
The addition of SE attention to the PC model (PC + SE) further enhances segmentation performance, particularly for smaller tumor regions, as evidenced by DSC improvements of for WT and for ET. However, a slight reduction in TC performance ( decrease in DSC) suggests that while attention mechanisms provide benefits in some areas, they may introduce trade-offs for certain tumor regions. Several lightweight attention modules are compared in Section 5.2.
@@
This improvement underscores the utility of the WDL in case of a segmentation region size imbalance. However, we notice that when adding the attention mechanism, the result of WT and TC decreased slightly in both DCS and HD95, which leads us to check whether attention is needed. Nonetheless, we chose to include SE because there is a 1.28% improvement in the ET segmentation result and enhancements in HD95 for TC and ET. This is further investigated in Section 5.4.
Figure 5 depicts qualitative comparisons between the various networks in segmenting the distinct tumor regions. The qualitative results demonstrate that our LATUP-Net which consists of PC and SE trained with WDL, outperformed all other models, consistent with the quantitative results in Table 5.1.
@@ 5.2 Comparison of Different Attention Mechanisms
In our approach so far, we have used the SE attention mechanism. In this section, we explore whether alternative attention mechanisms (see Section 2.3) may yield improved performance within the confines of our model. For a uniform evaluation, we incorporate all attention mechanisms after the parallel convolutions (PC). It is also important to note that the model is trained using Dice loss and the results presented here are derived from configurations that do not include the weighted Dice score loss (WDL). This choice was made to isolate the impact of the attention mechanisms themselves, allowing for a clear comparison of their effectiveness in the absence of the WDL’s influence.
@@
Table 5.2 summarizes our findings, detailing the performance metrics of residual SE, CBAM, ECA, and their combinations. While all mechanisms enhance the model’s accuracy, SE stands out as the most effective.
CBAM, despite its comprehensive design, might be limited by its complexity, which could result in computational overhead, potentially affecting efficiency, especially in real-time scenarios. On the other hand, while ECA is designed to be lightweight and computationally efficient, it may not capture the intricate inter-channel dependencies as effectively as SE. This trade-off between computational complexity and attention performance suggests that CBAM and ECA may be better suited for specific tasks or resource-constrained environments where accuracy can be sacrificed for speed or memory efficiency.
The SE mechanism’s advantage lies in its ability to efficiently highlight enhanced tumors, which is crucial for tasks involving small yet important regions like the tumor core (TC) and enhancing tumor (ET). Additionally, the reduced number of parameters in SE makes it particularly advantageous for large-scale medical image segmentation tasks, where model size and efficiency are important factors. The residual SE mechanism further boosts performance, particularly in terms of Dice similarity coefficient (DSC), as shown by its slightly higher values across all tumor regions.
Analyzing the standard deviation values, the variance in DSC scores appears minimal, which suggests that the choice of attention mechanism might not drastically affect the accuracy. However, in practice, a more streamlined model with fewer parameters (such as PC+SE) may offer competitive or even superior performance due to its balance of effectiveness and computational efficiency.
In contrast, the HD95 metric highlights more prominent differences across the methods, suggesting that attention mechanisms have a varying impact on capturing spatial discrepancies in segmentation, such as outliers or irregularly shaped tumors. For example, CBAM exhibits lower HD95 values for the whole tumor (WT), indicating that its spatial attention might better capture large-scale spatial variations, while SE-based models perform better on smaller, enhancing tumors (ET), where channel-wise attention is more effective at capturing fine-grained details.
Given the observed performance trade-offs, the choice of attention mechanism may depend on the specific task requirements, including segmentation accuracy, computational resources, and the need to balance spatial and channel-wise attention.
rowsep=3pt
{tblr}
colspec=Q[2cm,valign=m]lllllllll,
cell1-21-8=gray!20,
cell11=r=2l,
cell12=c=3l,
cell15=c=3l,
cell18=r=2l
Method DSC () HD95 (mm)
WT TC ET WT TC ET
PC+SE
PC+CBAM
PC+CBAM+SE
PC+ (SE-3D)
PC+Residual SE
PC+ECA
5.3 Lightweight Model Validation
To demonstrate the lightweight nature of LATUP-Net, Table 5.3 presents a comparison with other models on the floating-point operations (FLOPs) required to evaluate the model once, the number of parameters, and the inference times. Fewer FLOPs indicates a more efficient training process, as less computational power is required. Similarly, a lower number of trainable parameters not only reduces training time but also minimizes memory usage during deployment. Additionally, shorter inference times further highlight the lightweight characteristics of the model. It is important to note that there are various methods for estimating FLOPs. In our study, we used the TensorFlow profiler [52] function from the TensorFlow package, ensuring consistency and accuracy.
The baseline 3D U-Net, with million parameters and GFLOPs, is one of the more complex models, particularly in terms of parameters, reflecting its deeper and wider architecture. In contrast, Inception has fewer parameters ( million) but a much higher GFLOP count (), suggesting that while it has fewer parameters, its increased computational demand arises from its more intricate layer configurations.
Compared to Inception, LATUP-Net (PC + SE) achieves a more balanced and efficient design, with GFLOPs, million parameters, and an inference time of 168 ms, measured using an NVIDIA RTX 3060 GPU. It is important to note that inference and training times are highly dependent on the hardware used, which is why these times are not compared with other state-of-the-art models in this study. Despite this, LATUP-Net is far more efficient computationally. The PC-based variants (PC, PC + SE, and PC + Residual SE) maintain a similar parameter range ( – million) and low GFLOP counts ( – ), highlighting their efficiency. While models like PC + CBAM and PC + ECA slightly increase GFLOPs, they still maintain efficient inference times. Overall, LATUP-Net provides significant advantages over Inception by delivering better performance with fewer computations, making it a more efficient model, especially in terms of computational complexity and speed.
colspec=lccc,
cell11-4=gray!20
Models Parameters (M) FLOPs (G) Inference Time (ms)
U-Net
Inception
PC
PC + SE
PC + CBAM
PC + CBAM + SE
PC + (SE-3D)
PC + Residual SE
PC + ECA